antees.

AfterFinals is still desperately seeking a PC. I’m already weighted down with I-Banquet. The rest of my ExecComm is drowning in work. I—we— really need a brother to step forward and deal with this. It’s not that hard…

Don’t forget, current officer hours are:

• Dave Cho: Sundays, noon-1 pm

• Kevin Dahm: Thursday nights. I don’t remember when, but if you look at the office door, you can see when it is.

Come on by if you have questions. We like visitors, really we do. :)

Hey, Amy, how much more do I need to write? :)

And One More Dave

OK, Amy needs one more page. Not a problem. I can spew on for hours. This is called wasting toner. It’s also called an exercise in stream of consciousness. Or, in this case, stream of unconsciousness. Not that I’m really that tired; I just feel that tired. But it’s still kinda fun, at least for me, to blather on and on and on.

Another announcement! They Might Be Giants will be performing at Harvard University’s Alumni Hall on Thursday, April 28, 1994 at 9:00 pm. Tickets are $16.00, and are available from both the Holyoke Information Center and the Sanders Theatre Box Office.

I guess it’s not too early to talk about the summer now, especially since I’m most certainly looking forward to it. For starters, Nancy Clark and I (and possibly John Grossi) will be working on a Summer Sectional Fellowship Event. Or at least, that was the idea at last time; it might be a service project. So we’ll be needing help with it (surprise!), so if you’re around during the summer, keep your eyes open for this.

New topic: On July 4th, it’s traditional that a lot of people descend on the Esplanade to listen to the Boston Pops perform and watch Project Report: 1994 Institute Screw 23

The Preparation . . . . . . . . . . . . . . . . . . . . . . . . 23

The Voting and the Results . . . . . . . . . . . . . . . . . . 24

The Crew . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

CPPC Report 25

Stuff from mrn: Random News on AX Alums 29

In Memoriam: Earle M. Herbert 29

Chapter Meeting Summary: April 6, 1994 30

ExecComm Summary: April 13, 1994 32

More Spring Breaking 33

More Dave 34

And One More Dave 35

Advertisements 36 Harry, Ann, Joe, Barbara, Amy, and Ping, not necessarily in that order.

The SVP reported a past project (Easter eggs), a once-present-but-now-past project (Big Screw), a once-future-once-present-but-now-past project (Cub Scout Decathlon), and a future project (Afterfinals, which will be at Camp Resolute).

The MVP reported her hosedness.

The AVP reported that she was spending too much time running around trying to arrange stuff for the PMD party we’re sponsoring.

The treasurer had nothing to report.

The fellowship chair asked everyone to attend Steak Fry, to PC I-Banquet, and to go to The Gondoliers and/or Little Shop of Horrors.

The APOcEd reported that the APOc existed and distributed them.

The historian reported that he did the SFF page for the scrapbook.

The PubDir had nothing to report.

The ASC reported (from afar) that she was “at a test (ugh) review. but if [she] was there, [she]’d actually have things to say, like: i cleaned up the office last week, and then *poof* a couple of days later it was darn hard to tell. it shouldn’t work this way. PLEEEEEEEEEEEZE clean up after yourselves… ”

The Chairman of the Advisory Committee had nothing to report.

The Big Screw PC reported that he hoped to get 2.70 to run.

The President reported that this was the first chapter meeting in three weeks, that hopefully replies for sectional service project would be showing up soon, that we have a fridge and a stereo system in the office, that as a one time exception ExecComm last week agreed to sponsor a PMD (People Making a Difference) fundraising party in Walker, and that he was willing to co-PC Afterfinals with a pledge.

As for new business, the following legislation was proposed for submission to Nationals:

Moved: That Item 7 under “Procedures” (The Ritual and Other Ceremonies) be amended to read: living in e40 so I don’t get to the office more than once a day (this is a lot less than it used to be. :) Some days I don’t visit the office at all :)). I will be probably be taking Japanese III and a random management class while working principally on my research.

Reason for running now: I thought this term was great; I like the higher level of enthusiasm and fellowship that Dave has inspired in the chapter. Dave has/had plans for the chapter, some of which still need to be nursemaided a little. I can do this while implementing some of my own ideas or other people’s ideas. I get the whole summer to motivate myself and everyone else and to get ExecComm together and get organised. The chapter succeeded in its goal of getting a large pledge class, but now we are faced with the challenge of turning this into an increase in the number of active brothers.

Reason for not running now: A term off ExecComm would probably be good. For those of us who are on ExecComm far too often, we often forget that we see the chapter very differently from those brothers who are less involved in decision-making and planning. It always feels good to be “just a brother.”

Kate

President: Bridget Spitznagel

This is about the first time I’ve remembered to write a candidate article before the relevant elections meeting, so I’m really going to give the APOc its money’s worth. :-)

Hi. I’m Bridget and I’ll be running for president this evening :) For your amusement I may be running for some other offices afterward. I would wear my Australian hat again, but I lost it earlier this month (wah).

What do I think the president does?

The official duties of the prez involve various kinds of coordination. (Fortunately, no hand-eye coordination.) To make a silly paraphrase, each officer and committee within the chapter deals with a different subset of the things necessary to keep the chapter running in the present and to prepare it for the future; it’s the president’s job to make sure the separate parts are working together and to keep the

4 After some time, the pledges spoke up and mentioned how they weren’t getting the information they needed nor when they needed it. The rest of CPPC was devoted mainly (1:45) to Getting Clue To The Pledges; the last 20 minutes was about Making Pledges Part Of The APO Community. The other topics above were not discussed.

The part about Getting Clue To The Pledges mainly emphasized the point that the brothers are not telling the pledges neither what they need to know, nor when they need to know it. Pledge manuals are far too late; they should be coming out at Pledge Ceremonies, as usually planned. Pledges don’t necessarily know “what’s next,” what they should do next, how they can find out about it, etc. Too much that is automatically assumed by the brothers (the existence of signup sheets, what they’re for, and where they exist) is screwing things up.

Ideas/solutions:

• HTML [Editor’s Note for the Non-Nerdy: HyperText Markup Language] the pledge manual? 1/2 :)

• There are four varieties of contacting people: email, phone, snail mail, and personal. Know the advantages and disadvantages of each.

• The pledge manual needs to have very basic information. Things like:

– signup sheets exist

– email lists exist

– chapter meetings and ExecComms exist; and for this and the above, explaining what they are.

• A mandatory reception after the pledge ceremony is a good idea. During this, feed them ice cream, give them an office tour, give them their pledge manuals, and a pledge manual quickref.

• The quickref should explain very basic things in two ways: what, and how.

• A real big problem for pledges is: What now? What do I do now? Where should I go next? Who should I talk to? and was elevated in Fall 1989. Since that time, I’ve been variously involved with the chapter, most notably as a uniquely lame Fellowship Chair in Spring 1993 and now currently as President. I plan on graduating this December.

Next term, I’ll be taking 6.044 (? Al Drake’s algorithms course, which I hear isn’t terribly hard), two music courses (yay!), and thesis.

Why. Why do I want to be MVP? Well, a little history. Way back when I was a pledge (both terms), I said to myself, “Someday I want to be MVP and President.” No, not simultaneously. So this isn’t a whimsical thing. Nor is it because the slot is empty. Nor is it because I think the other candidates are lame (of course, I know of no other official candidates). It’s because I really want to be MVP. I’ve seen a lot of MVPs, and their different styles, and the problems they’ve addressed. I’ve worked on the pledge program by helping pledges formulate their problems; I’ve worked with specific brothers in identifying and addressing problems; I’ve worked with MVPs directly. But I haven’t done the job personally, and I would like to.

I’m old. I’ve seen things the way they’ve been done before. I understand the chapter, or at least, I think I do. :) I have confidence in myself, in my abilities, in my understanding of things.

I love this chapter, and the people in it, and what we do. That’s why I ran for President: because I was interested in doing more, in making the chapter grow, in making this chapter better. And a little secret plan I’ve had for awhile was to be Prez and “set things up” for a really keen term as MVP. So I’ve been trying to get things into order; to get brothers used to the idea of greatness. And the logical followup is to bring a pledge class to take up the work where I left off (oh God, this is really stupid, but i’m tired and it’s 12:37 am, so please bear with my inane writing). But it’s a dream, and I’m working on it, so there. :)

Now, before anyone gets a really stupid idea, of COURSE I’ll understand if you flush me like a dead fish. Don’t be silly. :) Besides, it just means I’ll have time to do Concert Choir next term, and I hear we’re doing something really cool, and you should come to our concert THIS FRIDAY AT 8 PM IN KRESGE; TIX ARE ONLY $2 IN ADVANCE SO LOOK FOR OUR BOOTH IN LOBBY 10.

8 tually understood most of it. Amazing. Napped. Discovered that Domino’s Extravaganza pizza on thin crust doesn’t give me too bad of a grease cough. Napped. Read some 12.001 (a. k. a. “Rocks for Jocks,” but the class is actually a pretty big time sink).

Tuesday: Didn’t sleep late, darn. Worked desk at New House, wrote three letters, delivered mail, gorged myself on Smartfood, sour cream and onion Pringles, and Chewy Chips Ahoy cookies, read more 12.001, napped, worked on the APOc, went to a surprise party for Emily Berkowitz of the Alumni Office (soon to be Emily Sandberg!), had cake, worked on the APOc, went to the Tech news meeting and somehow managed to pick up two stories to write for the next two issues, ate Chinese food, and came back to the APOc… where I still am (April 20, 12:57 am).

YiLFS, Amy

P. S. Final copy of this issue printed a little after 1:55 am. Whew.

Project Report: 1994 Institute Screw

The Preparation

The candidate search began Registration Day and ended the day the voting ended. Numerous potential candidates were approached, both faculty and staff. It is amazing how professionally these people can say no. “Foo is a very sensitive issue right now.” “I’m not that thick-skinned.” But these and other answers did not stop the quest for candidates. Publicity included an LSC slide, a drop poster, flyers, and word-of-mouth advertising. The dexion was set up Sunday, April 3, 1994 (the year for anyone reading this in the distant future :) ).

Treasurer: Henning Colsman-Freyberger

I am running for Treasurer.

1)

“class”  units
11.001J Intro to Urban Design and Development  3-0-9
11.011 The Art and Science of Negotiation  3-0-9
14.24 Economics and Law  3-0-9
14.54 International Trade  4-0-8
Fencing Team  8-2-0
German House Steward  0-2-2
New House ComCom Chairperson (pending election) 0-3-1
New House Vice President (pending election)  .5-2-1
SafeWalk  0-2-0

Total: 21.5-11-39 = 71.5 units

2) a) I have undergone Treasury training and feel I can competently carry out the duties of this office.

2) b) I have undergone Treasury training and feel I can competently carry out the duties of this office.

Henning Colsman-Freyberger

APOcrypha Editor: Bert Dvornik

I haven’t quite decided whether to run for APOcEd or not. However, I’m going to write and submit an article. I can always decide to back out before [or during] elections.

I have never done any newsletter editing before, but I like to play with LaTEX, and I believe I can learn the tricks of the trade that I don’t know already. Anyway, editing the APOcrypha seems like a fun job.

I’m going to make only one campaign promise: I’m going to be getting APOcs out a day or two before the chapter meeting. Thus, the brothers will be able to read anything relevant to the meeting before the meeting itself.

For Ezra’s benefit, I’ll admit that I don’t really know what classes

12

A Random Article…

Hiya…

I’m just a brother who’s gone incommunicado for a while, all because of classes. :-( I even had to put off volunteering at the hospitals and forego fun chapter projects to keep from flunking a few classes. Yup, I do know a few of the pledges on the list… but I haven’t had a chance to talk to them. Even Mike, whom I’d met during Rush last year. I do hang around the APO office when I get a chance (usually at rather odd hours) … and just recently, I saw how Ping printed some cards on that odd printing press. I’ve resolved to learn printing the very next chance that I get. I even read a children’s book on Hiroshima that was given to APO (I recommend it to everyone… it’s on the APO bookshelf near the historian’s desk). Has anyone compiled the top ten reasons to go to MIT yet? Or better, the top ten reasons for being in APO?

Well anyways, hope to see many of you around (and I’ll try my best to “reactivate” myself). Until then, life is just one indefinitely long night of tooling.

YiLFS, Cindy! :-)

The office got broken into

The following report was written by Kevin; I’ve added a few comments.

You have probably heard some references to the incident involving missing money from the Big Screw contest. This is to update you on everything we currently know about the situation.

Thursday night, after the Big Screw money for that day was counted, Henning (Big Screw PC) placed it in the office strong box, locked it, and returned the key to its place in the treasurer’s desk. Friday around lunchtime, the treasurer’s desk drawer was seen wide open. Friday evening when Henning went to retrieve the money from the strong box, all the bills were missing. It was subsequently determined that a few dollars were also missing from the treasurer’s desk.

Thus, the most likely scenario is that late Thursday night, someone gained access to the APO office, took the money and key from the treasurer’s desk, opened the strong box and took the money, locked the strong box, returned the key to the treasurer’s drawer, and left.

Any additional information you may have should be brought to my attention immediately.

WE INTEND TO TAKE THE FOLLOWING ACTION:

1. The chapter will reimburse the Big Screw contest out of the general service fund. (Recommended at last Wednesday’s ExecComm.)

2. Effective immediately, cash will NOT be left in the treasurer’s desk drawer. It should be placed in an envelope with any accompanying receipts or messages and slid into the strong box, which will be kept locked at all times.

3. The key to the strong box will no longer be kept in the treasurer’s desk, the president’s desk or any other accessible location in the office.

In addition, I remind you that the office door must be kept locked at all times unless someone is present in the office, and that someone

16 been kicking around for a while (such as a quotes section, or a 25 Years Ago in the APOcrypha section, an idea with which readers of the Scientific American magazine will be familiar).

YiLFS, Ping Huang

PTRM!

This is a reminder that Pledge Term Review Meeting (PTRM) is

THIS SATURDAY, APRIL 23

The pledges (once again) are:

Shiu Au  Richard Barbalace Angela Chang
Helen Chen Cathy Coury Andrew Garcia
Hannah Jang Aphrodite Kakouros Juliana Leung
Cris Lin Jeffrey Poore Brad Ricketson
Mike Saginaw Andy Su  Grant Su
Junko Tamiya Emily Wallis  Bryan Watabe

President’s Corner

Hiya. I was lame last time and didn’t write for the APOc. Sorry, Amy. Nor have I written a “what I did during Spring Break.” Soon. :)

A few noteworthy events happened over the last four weeks: 1, Big Screw, 2, the office got broken into, and 3, CPPC.

Big Screw

Big Screw happened. Good work, Henning, and everyone involved. I think perhaps next year, we really do want to have a contest ¿ five days. It’s just too hard for the professors to campaign otherwise; the profs really need to mention it in lecture.

Henning’s Random Ramblings III

The reason I asked last APOc’s trivia question (What is the color of the covers of United States passports currently issued?) is that it served as a reminder to renew mine. This I have done in the meantime, so I am well informed now. The correct answer is dark blue. Congratulations are in order for Ann and mrn, who got it right. I had thought that the color had been changed recently and so I was surprised to be informed that my passport would be dark blue again. After further inquiry, my suspicions were confirmed. The color of the passport had changed about 18 months ago to an olive drab, but had been changed back to the blue a year later.

This week I thought I’d pose three questions:

• One to encourage office exploration: What are the flavors of the sodas sold by the APO Coke machine and what is the price?

• One to promote exploration of Athena: What is the ratio of “No Smoking” to “No Food or Drinks Allowed” signs in the student center cluster?

• One to challenge non-MIT intellect: What is the political status of Greenland?

That should provide enough scope. :) As usual, send your answers to trivia@mit.edu. Also, I am offering a new service. If you have any questions about anything APO-related, feel free to ask them and I will answer them in the next issue of the APOc. Anonymity is guaranteed if requested.

Henning Colsman-Freyberger

Pledge Talk

Well, it is almost the end of the term. A lot of people are getting hosed now, and finals are just around the corner. But as everyone knows, so is PTRM. And soon is the meeting where pledges will be initiated into brothers. I figure that I must restate what I said at

20

Administrative Vice President: Gilbert Leung

I somehow have succumbed to the charm of Amy and started writing this article that you are reading. What is this article about again? Hmm… right! I’m running for AVP! I’m psyched! I’ll have time! Vote for me!

What? You are not convinced? Ok. I’ll tell you a reason or two why I want to run.

1. I want my name added to apo-acl, the real power! (In case you don’t know, apo-acl is the athena list that controls all the apo mailing lists as well as the apo locker. Once you are on the list, you will have the permission to nuke everything in the locker. >:) ok. ok. That’s not exactly what I’m going to do…)

What? That is still not convincing enough? Hey, when I said a reason or two I meant it! So did I when I said that I’m psyched. Hmm… Voters are certainly more and more demanding everyday… ok. ok. Let me make up, err… tell you another reason:

2. I never really pay 100% attention at meetings (which I consider bad). If I become the AVP, I will have to take minutes and will force myself to pay heed (which I consider good). Or, even better, I shall learn how to pay only 50% attention and still manage to take minutes. :)

So, I have cited my two reasons and don’t you make me contradict myself by asking me to give you another!

Gilbert

The Voting and the Results

Votes ($) were cast 10 am-5 pm each day, April 4-8. The final tally:

Robert Logcher 1.00 $179.12
Jacob White  6.003 $138.00
Irwin Oppenheim 5.60  $56.74
Stephen Lippard 5.03  $44.75
John Guttag  6.170 $30.30
Bill Clinton  Hack  $4.96

A total of $492.80 was raised. It will go to the Planned Parenthood League of Massachusetts. New policies concerning the handling of money resulted from the contest [see ExecComm Summary].

The Crew

Candidates PC  Jerry Lin
Manpower PC  Bryan Watabe
Money PC  Shiu Au
PC  Henning Colsman-Freyberger
Publicity PC  Richard Barbalace
Publicity PC  Dima Dribinsky

Many thanks go to the PCs and the many workers who manned the booth. Without all the help, this logistical task wouldn’t have been the success this year’s Institute Screw Contest turned out to be.

YiLFS &

Henning Colsman-Freyberger

1994 Institute Screw Project Chair

24 and pledge ceremonies, provided they are planned at least a week in advance.)

What else?

If not elected president, I may run for APOcEd (I forget who’s running already.) I know LaTEX, and I would put out APOcryphae regularly (and probably even remember who to mail them to :) ). I can’t guarantee content (you can only twist arms so far before they break off and their owners can’t type anymore), but if I didn’t get any actual content I might stick in some useless but amusing things, such as a pack of lies column, or a how to cook unusual things in a microwave column; or I might reprint interesting/useful bits of the pledge manual, the chapter book, how-to documents, etc.

If inadvertently elected Fellowship Chair, I promise to be interesting (you can take that as a promise or a threat; either way).

YiLFS, bridget

Membership Vice President: Dave Cho

I’m running for MVP. This article addresses the following topics in the following order:

1. Who am I? A brief introduction.

2. Why

• Why do I want to be MVP?

• Why I shouldn’t be MVP.

3. What would I do?

4. How would I accomplish my goals?

5. When will I have accomplished these goals by?

6. Where would I hold events?

Who. My name, just in case any of you forgot, is dave cho. I’m currently a first-term senior in VI-3. I first pledged in Spring 1989 • The tradition of Big Brothers pinning their Little Brothers at

Pledge Ceremonies should be reinstated. This will require a pre-Pledge Ceremony meeting, which is a good idea anyways to explain more of the chapter.

• Information density, especially of the office, is a problem. The pledges are being hit with a firehose, and don’t have a context to process it with.

• Everything in the office should be labelled, e.g., “This is the Signup Sheet board. Signup sheets, except when at meetings, always live here, so you can come in any time and sign up for project.”

• Open Meetings need to be restructured again; prospectives still aren’t getting the information they’re interested in. For instance, we should give them a “roadmap” for the future: what they need to do, what meetings they need to come to, how to find out more, etc.

It should be stressed that the current inadequacy with the pledge program is the result of many years of problems. Some of the above problems stem from recent sliding of discipline over the last few terms; some have been prevalent since I was a pledge six years ago.

Onwards.

We finished by discussing the problem of incorporating pledges into the “APO Community.” Suggestions included

• We should explicitly point out that brothers aren’t trying to find reasons to not elevate pledges, but rather, reasons to elevate.

• A concern of many brothers is legacy; in a few years, we’ll be turning over the chapter to these pledges, and we want to make sure that it’ll be in good hands.

• Personally ask for a pledge’s opinion. Telling a group of pledges we want to hear from them does not have the same effect as going up to a pledge and specifically asking them for advice, help, assistance, etc.

28 during ceremonies) that we are proposing? The chapter maintains a National Layaway fund that can be used to defray the expenses of going to Nationals, so don’t just dismiss the idea out of hand. (On a personal note, seeing as how Nationals is between Christmas and New Year, I will be in Dublin and not in Dallas).

Meetings: Should always be short, interesting and well-run. I can try. :) People seem to like knowing that meetings will not normally run over an hour. I like meetings where something worthwhile is accomplished; otherwise, spending even an hour on it seems too much. I would also like to differentiate better between chapter meetings and ExecComm meetings, although right now I’m not sure what belongs where… Brothers should be equally welcome at either, but they are intended to serve different needs.

ExecComm: Should work together to make the term a success for all the brothers and pledges and alumni and other people. Regardless of responsibilities of the individual office, each person’s general contribution to ExecComm is very important. I’d also like to see lots of new faces, together with the old (not just agewise) ones.

By-laws: As a personal goal I would like to get together a small group of people interested in amending the by-laws so that they are consistent and practical. For example, the references to the budget are inconsistent and the by-laws have no escape clauses for when you try to elect someone for an office for which there are no candidates. The recent national by-laws discussions also raised the question about inserting an anti-discriminatory statement in our by-laws, since it was removed from the national by-laws. I am sure that many more issues will come to light when we start discussing them.

Alumni: They’re nice people, but they’re sometimes a little confused about what the chapter needs from them, or wants from them, or thinks it needs or wants, or just thinks in general. Do we want a Big Uncle sort of program? Do we care? Why did the idea not get implemented anyhow? Were we lame? Were they lame? Were we all lame? (Probably) I can try and talk to the alumni association to see what’s happening.

Finally…

In answer to Ezra’s favourite question: I haven’t a clue what I will be taking next term. I’m a grad student. I have an RA. I will be Use of blindfolds is optional. Where chapters choose not to use blindfolds, or where state or local law or institutional regulations prohibit blindfolding, substitute closed eyes where appropriate.

Rationale: We feel that the use of physical blindfolds does not necessarily contribute anything to the effect of the ceremony which cannot be achieved equally well by simply having pledges close their eyes. Further, use of blindfolds is awkward and can be considered intimidating, while our asking the pledges (or neophytes) to close their eyes shows that we trust them.

The motion to submit this amendment to Nationals was carried many-0-1 by voice vote.

As for DSKs, the following people were nominated and discussed: Ezra Erb, Dave Cho, Kate Curtis, Bridget Spitznagel, Barbara Manganis, Crystal Reul, and Mike Person.

PFTGOTOs were as follows:

Amy: Thanks for making this issue of the APOc the biggest yet! Yale: Congrats to Kate on getting the fridge! Dave: Don’t send non-APO-specific email to apo-news. Henning: Thanks to 6.003, 1.00 and 5.60 for Big Screw.

ExecComm Summary: April 13, 1994

Start time was 7:32 pm. People were there. Reports that were not time-sensitive are as follows:

Historian (Dima): Still working on matching names to pictures.

Advisor (Joe): I have the certificates for everyone who took the CPR class.

Silkscreen (Dima): Rachel wants to put together a schedule of what runs people are planning to do.

President (Dave): I will be PCing I-Banquet. It will be at 100 Memorial Drive. Initiation will be on the third floor of building 26.

New business: For a full report on the facts of this business, please see Kevin or Dave who have written up a report.

32

Contents

Editor’s Sigh of Relief 1

Candidate Articles! 1

President: Kate Curtis . . . . . . . . . . . . . . . . . . . . . 1

President: Bridget Spitznagel . . . . . . . . . . . . . . . . . 4

Membership Vice President: Dave Cho . . . . . . . . . . . . 7

Administrative Vice President: Gilbert Leung . . . . . . . . 11

Treasurer: Henning Colsman-Freyberger . . . . . . . . . . . 12

APOcrypha Editor: Bert Dvornik . . . . . . . . . . . . . . . 12

APOcrypha Editor: Ping Huang . . . . . . . . . . . . . . . 13

PTRM! 15

President’s Corner 15

What I did for spring break… 17

Just another article. :) 17

Pledge Article: EGW 18

A Random Article… 19

Henning’s Random Ramblings III 20

Pledge Talk 20

Silly Stuff 22

Patriots Day Weekend: The Editor’s Journal 22 pretty fireworks. Assassins, SIPBions, and APOers convene in a traditional spot; watch your email for announcements at a later date.

The Minuteman Bike Trail connects between Alewife T Station and Belmont. However, there is another bike trail connecting Alewife and Porter. Therefore, if you bike down Hampshire street, you can, in relatively high safety, bike the fifteen miles between MIT and Belmont. On the way, you travel through Arlington and Lexington. People are already getting their bikes ready and starting to ramp up for summer. If you’re interested in joining this amoebic mob, send me mail.

Diskless is a group of people who play Ultimate Frisbee very poorly. I mean, we suck. Correction: we really suck. As member Chad Brown describes it, “We often know where we want the frisbee to go.” Currently, some subset of these people meets in Killian Court at 5:30 pm on Fridays to play. Come on by; if you suck, you’ll fit in perfectly. Or you can cheer; a lot of people just spectate.

OK, I’m afraid this’ll be too long, so let me wrap it up.

dave

Advertisements

WANTED: APOcrypha submissions of all shapes and sizes for the final issue of the term, premiering at I-Banquet, May 6. Must be two-dimensional and photocopiable. Titles will be awarded to distinguished contributors! Send your submission through email to

aihsu@mit.edu or apocrypha@mit.edu (whichever you feel like typ-

ing) or leave it in the APOcEd’s box in the APO office, or address it to Amy and pin it to the bulletin board, or address it to Amy and pin it on the office door, or address it to SPOO and leave it in the Technique office, or send it interdepartmental mail to Amy Hsu, New House, or send it U. S. mail to 476 Memorial Drive, Cambridge, MA 02139, or wave it in my face and say, “It’s for the APOc! Lose it and die!” or… well, you get the picture. Submit!

STILL HUH? Poor APOcEd is still having problems with page numbers. I LaTEXed the silly thing up to ten times and it didn’t seem to work. HELP!